Which of these expressions is Little's Law?

Inventory/Throughput Rate = Flow time

Throughput Rate/Inventory = 1/Flow Time

Inventory/Flow time = Throughput rate:

Inventory = Throughput rate × Flow time

All of the above
Answer #1

1. D. All of the above

Little’s law represents the following equation:

RT = I, where R = throughput, T is the flow time and I is the inventory. Using the equation, we can derive the same answer using different variations in the formula, which equates to the same value. Therefore,

R * T = I can also be written as I / R = T, R / I = 1 / T and I / T = R